If [x] is the greatest integer not greater than x, then determine the limit of [x ]when x goes to (1/2) is it exists


1
Expert's answer
2021-02-02T05:23:52-0500

Here, we have "f(x)=[x]" (greatest integer function)


The graph of this function looks like this:-




Here, we want to find "\lim\limits_{x\to\frac{1}{2}} [x]"


Here we calculate LHL and RHL separately.


LHL:-


"\lim\limits_{x\to\frac{1}{2}^-} [x]=0" (As here x approaches "\frac{1}{2}" from left and so, [x] is 0)


RHL:-


"\lim\limits_{x\to\frac{1}{2}^+} [x]=0" (As here x approaches "\frac{1}{2}" from right and so, [x] is 0)


"\therefore" As LHL=RHL=0 , we know the limit exists and is:-



"\color{purple} \lim\limits_{x\to\frac{1}{2}} [x]=0"